OpenStudy (nuccioreggie):

Simplify by dividing negative three over eight divided by five over nine.. negative fifteen over seventy-two. fifteen over seventy-two. negative twenty-seven over forty. twenty-seven over forty.

OpenStudy (texaschic101):

(-3/8) / (5/9) when dividing fractions, you " flip " what you are dividing by, then multiply. Make sure you flip the right number. (-3/8) * (9/5) --- see how I flipped what I was dividing by...now just multiply -27/40 and thats it :)

OpenStudy (whpalmer4):

Yep, just as I explained in your previous question. \[\frac{\frac{a}{b}}{\frac{c}{d}} = \frac{a}{b}*\frac{d}{c} = \frac{a*d}{b*c}\]

OpenStudy (moazzam07):

so i the answer -27/40 @texaschic101